OpenStudy (anonymous):

whats the slope-intercept form perpendicular to y=1/2x-1, x-intercept at 4

OpenStudy (anonymous):

\[y - 0 = -2(x - 4)\]\[y = -2x + 8\]

OpenStudy (saifoo.khan):

perpendicular slope = \[m_2 \times m_1 = -1\] \[\frac12 \times m_1 = -1\] \[m_1 = -2\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

saifoo can you help me out?

OpenStudy (saifoo.khan):

i read your question. but i was unable to get it.. :?

OpenStudy (saifoo.khan):

:(

OpenStudy (anonymous):

its + 8 saifoo not + 4 :)

OpenStudy (saifoo.khan):

LOl, i read that as "y-intercept" . my bad.

OpenStudy (anonymous):

lol no problem :D